The following are the cash flows of two projects:

Year Project A Project B
0 $ (260 ) $ (260 )
1 140 160
2 140 160
3 140 160
4 140

What are the internal rates of return on projects A and B? (Enter your answers as a percent rounded to 2 decimal places.)
Answer #1

To find IRR we need to use trial and error method. We need to find two discounting rates so that one when we discount total cash inflows should be greater than cash outflow and other should be less than cash outflow. You can choose any two rates, but rates which are closer to IRR gives better results.

Year Project A Discount Factor @ Discounted Cash Discount Factor @ Cash flows 38% flow at 38% 40% 140 0.7246 101.45 0.7143 14

IRR for project A

266.84-260 = 38+ 266,84-25g go * (40 - 38)

- 38 + (0.86038+2)

- 38 +1.206

- 39.7%

Therefore IRR for Project A = 39.72%

Project B:

Year 37% Project B Discount Factor @ Discounted Cash Discount Factor @ Cash flows | flow at 37% 39% 160 0.7299 116.79 0.7194

IRR for Project B =

264.26 - 260 = 37 + - 264.26 257.5 * ၄E * (39 - 37)

- 37+ (0.63 +2)

- 37+1.26

- 38.26

Therefore, IRR for project B = 38.26%

Formula for discount factor:

Discount Factor = 7 (1 + i)

Where,

i = discount rate

n = number of periods

In my calculations i have not rounded off intermediate calculations.